### v1

# Minimal Degrees, Volume Growth, and Curvature Decay on Complete Kähler Manifolds

Abstract.
We consider noncompact complete Kähler manifolds with nonnegative bisectional curvature. Our main results are: 1. Precise estimates among refined minimal degree of polynomial growth holomorphic functions and holomorphic volume forms, $\operatorname{AVR}$ (asymptotic volume ratio) and $\operatorname{ASCD}$ (average of scalar curvature decay) are established. 2. The Lyapunov asymptotic behavior of the Kähler-Ricci flow can be described in terms of polynomial growth holomorphic functions. This provides a unifying perspective that bridges the two distinct proofs of Yau’s uniformization conjecture. These resolve two conjectures made by Yang.
1. Introduction
As part of Yau’s program to study complex manifolds of parabolic type, he proposed the following well-known uniformization conjecture in 1970s:
Conjecture 1.1 (Yau’s Uniformization conjecture, [29]).
A complete noncompact Kähler manifold $(M^{n},g)$ with positive bisectional curvature is biholomorphic to $\mathbb{C}^{n}$.
In the maximal volume growth case, Liu’s breakthrough [16] confirmed the conjecture by combining Gromov-Hausdorff convergence techniques with the three-circle theorem developed in [13]. An alternative proof, based on the Kähler-Ricci flow, was later provided by Lee-Tam [8], building on results in Chau-Tam [3].
These advances led to a deeper understanding of the structure of such manifolds. Notably, through the works of [20, 21, 24, 11, 12, 17], the following conjecture of Ni has been established as a theorem:
Theorem 1.2 (Corollary 3.2 in [20], Theorem 1.2 in [24], Theorem 2 in [12], Theorem 1.4 in [11], Corollary 2.16 in [17]).
Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature. Assume that the universal cover of $M$ does not split. Then the following conditions are equivalent:
(1) $M$ is of maximal volume growth, i.e.
$\operatorname{AVR}(M,g)=\lim_{r\to\infty}\frac{\operatorname{Vol}\left(B(p,r)\right)}{\omega_{2n}r^{2n}}=\nu>0.$
Here $\omega_{2n}$ is the volume of the unit ball in $\mathbb{C}^{n}$.
(2) There exists a nonconstant polynomial growth holomorphic function, i.e. $\mathcal{O}_{P}(M)\neq\mathbb{C}$.
(3) The average scalar curvature decay is finite, i.e.
$\operatorname{ASCD}(M,g)=\limsup\limits_{r\to\infty}r^{2}\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p,r)}S$
is finite. Here $S$ is the scalar curvature.
Remark 1.3.
(1) Note that both $\operatorname{AVR}(M,g)$ and $\operatorname{ASCD}(M,g)$ are independent of the choice of $p$.
(2) A very recent result of Liu [18] implies that for any $(M,g)$ with nonnegative bisectional curvature, the $\limsup$” in the definition of $\operatorname{ASCD}(M,g)$ above can be replaced by $\lim$”, see also [22].
In his thesis [27], Yang defined:
Definition 1.4.
Let $(M^{n},g)$ be a complete Kähler manifold with nonnegative bisectional curvature, for a fixed $p\in M$, define
[…]
where $\mathcal{K}_{M}$ is the canonical line bundle on $M$, $\mathcal{O}_{P}(M)$ is the space of all polynomial growth holomorphic functions on $M$ and $H^{0}_{P}(M,\mathcal{K}_{M})$ is the space of all polynomial growth holomprhic $n$-forms on $M$. Set $d_{\operatorname{min}}=+\infty$ or $D_{\operatorname{min}}=+\infty$ if $M$ does not admit any nonconstant holomorphic functions of polynomial growth or $\mathcal{K}_{M}$ admits no nonzero holomorphic section of polynomial growth.
Now assume $(M^{n},\omega)$ satisfies the assumption in Theorem 1.2 and of maximal volume growth. Since the Kodaira dimension $K(M^{n})=n$, $\mathcal{O}_{P}(M)$ is “holomorphically regular” in the sense that we can always find “local coordinate by global polynomial growth functions”. The refined minimal degree can be also defined in [27] by
$\overrightarrow{d_{\operatorname{min}}}(p):=\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right):=\inf_{\{f_{1},\cdots,f_{n}\}}\left\{\limsup_{x\rightarrow\infty}\frac{\log|f_{1}(x)|}{\log d(x,p)},\cdots,\limsup_{x\rightarrow\infty}\frac{\log|f_{n}(x)|}{\log d(x,p)}\right\}$
where the infimum is taken among any $n$-tuple of global holomorphic functions that gives local coordinates at $p$ with the corresponding $\limsup_{x\rightarrow\infty}\frac{\log|f_{i}(x)|}{\log d(x,p)}$ arranged in a non-decreasing order for $1\leq i\leq n$, we denote this set by $\mathcal{O}_{P,p}(M)$. In other words, for any $1\leq k\leq n$ we have $d_{\operatorname{min}}^{(k)}=\inf_{f_{k}}\limsup_{x\to\infty}\frac{\log|f_{k}(x)|}{\log d(x,p)}$, where the infimum is taken among all possible $f_{k}$ that appears in the $k$-th component of some sequence in $\mathcal{O}_{P,p}(M)$. Note that apriori it’s unclear if $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$ can be obtained by an n-tuple holomorphic functions in $\mathcal{O}_{P,p}(M)$. But obviously $d_{\operatorname{min}}=d_{\operatorname{min}}^{(1)}$.
Then Yang proposed the following conjecture in [27] on the relation between the above quantities which can be understood as the quantitative version of Theorem 1.2.
Conjecture 1.5 (Conjecture 2.5.6, Conjecture 2.5.8 in [27]).
Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature. Assume the universal cover of $M$ does not split. Then
(1) $\overrightarrow{d_{\operatorname{min}}}(p)$ can be realized by a n-tuple holomorphic functions in $\mathcal{O}_{P,p}(M)$ and is independent of the choice of $p$.
(2) $\operatorname{AVR}(M,g)=\prod\limits_{i=1}^{n}\frac{1}{d_{\operatorname{min}}^{(i)}}$.
(3) $D_{\operatorname{min}}=\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n$.
(4) $\operatorname{ASCD}(M,g)=4nD_{\operatorname{min}}$.
In the case of nonmaximal volume growth, (1),(2) are true because $d_{\min}=+\infty$ from Theorem 1.2, and $\operatorname{ASCD}=+\infty$.
Assume $M$ is of maximal volume growth, in [27], Yang verified the conjecture is true for $U(n)$-invariant Kähler metrics on $\mathbb{C}^{n}$. Liu proved (1) implicitly in [16]. In fact, he showed that any n-tuple of polynomial growth functions which are algebraically independent and of minimal degrees can form a global coordinate. Then recently in [18], Liu proved (2) and derived the explicit formula $\operatorname{ASCD}(M,g)=4n(\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n)$ by passing the geometric quantities to the tangent cones and exploiting the metric Kähler cone structure of the limit space. In [28], Yang proved that (3) holds for gradient expanding Kähler-Ricci solitons with nonnegative Ricci curvature by using the Poincaré coordinate introduced in [1], see Theorem 1.8.
Therefore, the general proof of (3) remained the final step toward resolving Conjecture 1.5 in full generality.
In this paper, we prove that (3) of Conjecture 1.5 is true, thereby completely resolving the conjecture. Precisely, we proved
Theorem A.
Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature. Suppose the universal cover of $M$ does not split. Then
$D_{\operatorname{min}}=\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n.$
Now we move to the behavior of the Kähler-Ricci flow on complete noncompact Kähler manifolds where $(M,g)$ has nonnegative bisectional curvature.
$\frac{\partial g(t)}{\partial t}=-\operatorname{Ric}(g(t)),\quad g(0)=g.$ (1.1)
When the initial metric $(M,g)$ has bounded curvature and Euclidean volume growth, the long-time behavior of (1.1) has been described very clearly. Let us summarize these results in the following:
Theorem 1.6 (Theorem 1.2, Proposition 3.2 in [3], Theorem 6.1 in [4]).
Suppose that $(M^{n},J,g)$ is a complete noncompact Kähler manifold, where $g$ has bounded nonnegative bisectional curvature and Euclidean volume growth with $\operatorname{AVR}(M,g)=\nu>0$. Then the following conclusions hold:
(1) The solution to the Kähler-Ricci flow with the initial metric $g$ exists for all $t\in[0,+\infty)$ and it has nonnegative bisectional curvature for any $t\geq 0$. Moreover, there exists a constant $C(n,\nu)$ such that $\|\operatorname{Rm}(x,t)\|\leq\frac{C(n,\nu)}{t}$ for any $x\in M$ and $t\in(0,+\infty)$.
(2) For any point $x\in M$, let $\{\lambda_{1}(x,t),\cdots,\lambda_{n}(x,t)\}$ denote the eigenvalues of Ricci curvature $\operatorname{Ric}(x,t)$ with respect to $g(t)$ in the nondecreasing order. Then $t\lambda_{i}(x,t)$ is nondecreasing on $t>0$, hence $\mu_{i}(x):=\lim_{t\rightarrow+\infty}t\lambda_{i}(x,t)$ exists.
(3) If $\mu_{1}(x)<\mu_{2}(x)<\cdots<\mu_{l}(x)$ are the distinct limits in (2), where $l\leq n$. Then $V=T_{x}^{(1,0)}(M)$ can be decomposed orthogonally with respect to $g$ as $V_{1}\oplus\cdots\oplus V_{l}$ so that
if $v$ is a nonzero vector in $V_{i}$ for some $1\leq i\leq l$, and let $v(t)=\frac{v}{|v|_{g(t)}}$, then
$\lim\limits_{t\to\infty}t\operatorname{Ric}(v(t),\overline{v(t)})=\mu_{i}$
and thus
$\lim\limits_{t\to\infty}\log\frac{|v|^{2}_{g(t)}}{|v|^{2}_{g}}=-\mu_{i}.$
Moreover, both convergences are uniform over all $v\in V_{i}\backslash\{0\}$.
$\sum\limits_{i=1}^{l}-\mu_{i}(x)\operatorname{dim}_{\mathbb{C}}V_{i}=\lim\limits_{t\to\infty}\log\frac{\det(g_{i\bar{j}}(x,t))}{\det(g_{i\bar{j}}(x,0))}.$ (1.2)
(5) Fix any point $p\in M$. Given any $t_{k}\rightarrow+\infty$, define $g_{k}(t)=\frac{1}{t_{k}}g(t_{k}t)$. The pointed sequence $(M^{n},J,g_{k}(t),p)$ sub-sequentially converges to a gradient expanding Kähler-Ricci soliton $(N,J_{\infty},h(t),O)$ where $t\in(0,+\infty)$ in the following sense:
  1. (i)
    After picking a subsequence still denoted by $t_{k}$, there exists an increasing sequence of open subsets $O\in U_{k}$, which exhausts $N$ and a family of diffeomorphisms $F_{k}:U_{k}\rightarrow F_{k}(U_{k})\subset M$ with $F_{k}(O)=p$.
  2. (ii)
    As $t_{k}\rightarrow\infty$, the sequence $(U_{k},F_{k}^{*}J,F_{k}^{*}(\frac{1}{t_{k}}g(t_{k}t)),p)$ converges smoothly to another sequence of complete Kähler manifolds $(N,J_{\infty},h(t),O)$ uniformly on compact sets of $N\times(0,\infty)$.
  3. (iii)
    $(N,J_{\infty},h(t))$ has nonnegative bisectional curvature for any $t>0$, and there exists a real-valued function $f\in C^{\infty}(N)$ such that $(N,J_{\infty},h(1))$ satisfies the expanding soliton equation
    $\operatorname{Ric}_{i\bar{j}}(h(1))+h_{i\bar{j}}(1)-f_{i\bar{j}}=0,\quad f_{ij}=f_{\bar{i}\bar{j}}=0.$ (1.3)
    Moreover, $\nabla_{h(1)}f(O)=0$ and the eigenvalues of Ricci curvature of $h(1)$ at $O$, arranged in the non-decreasing order, equal $\mu_{i}(p)$ for $1\leq i\leq n$.
Indeed, in (3) in Theorem 1.6, $\mu_{i}$ play the role of Lyapunov exponents in a dynamical-system interpretation on asymptotics of (1.1), it basically says that the Ricci curvature can be “simultaneously diagonalized” near $t=+\infty$ in some sense. See Theorem 4.1 in [3] for a precise statement.
Then Yang proposed the following conjecture in [27]:
Conjecture 1.7 (Conjecture 2.5.16 in [27]).
Let $(M^{n},g)$ be a complete noncompact Kähler manifold with bounded nonnegative curvature and Euclidean volume growth. Let $g(t)$, $t\in[0,+\infty)$, be the complete solution Kähler-Ricci flow with initial metric $g$, then
(1) $\mu_{i}(p)$ is independent of the choice of $p$ for any $1\leq i\leq n$.
(2) $\mu_{i}=d_{\operatorname{min}}^{(i)}-1$ for any $1\leq i\leq n$.
(3) $\operatorname{ASCD}(M,g(t))$ is invariant along $g(t)$.
In [28], Yang proved (1) using Shi’s curvature estimates. He further proved (2) for expanding Kähler-Ricci solitons with nonnegative Ricci curvature. Precisely,
Theorem 1.8 (Theorem 1.4 in [28]).
Let $(N^{n},J,O,g,f)$ be a complete noncompact gradient Kähler-Ricci soliton with nonnegative Ricci curvature, normalized so that $f=R+|\nabla^{g}f|^{2}$. Let $\mu_{1}\leq\cdots\leq\mu_{n}$ be the eigenvalues of Ricci curvature at $O$. Then
$\overrightarrow{d_{\operatorname{min}}}(q)=\{\mu_{1}+1,\cdots,\mu_{n}+1\}\text{ for all }q\in N;$ (1.4)
and
$D_{\operatorname{min}}=\sum\limits_{i=1}^{n}\mu_{i}.$ (1.5)
If $(M,g)$ has unbounded curvature, in [8], Lee-Tam established the following long-time existence result for the Kähler-Ricci flow:
Theorem 1.9 (Corollary 1 in [21], Theorem 1.5 in [8]).
Suppose $(M^{n},g)$ is a complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth, $\operatorname{AVR}(M,g)=\nu>0$. Then there exists $C=C(n,\nu)>0$ depending only on $n,\nu$ such that there exists a smooth Kähler-Ricci flow solution $g(t)$ on $[0,+\infty)$ such that
$\|\operatorname{Rm}\|(g(t))\leq\frac{C(n,\nu)}{t}$
on $M\times(0,+\infty)$.
(2) $g(t)$ has nonnegative bisectional curvature.
(3) $\operatorname{AVR}(g(t))=\operatorname{AVR}(g(0))=\nu$ on $M\times(0,+\infty)$.
We emphasize that the uniqueness of such Kähler-Ricci flow solution remains unknown in this setting. Consequently, though these $\mu_{i}$’s can be defined for $M$ with unbounded curvature, a priori, we cannot guarantee that they are independent of the particular flow solution g(t).
In this paper, we prove the Conjecture 1.7 holds in general.
Theorem B.
Conjecture 1.7 holds.
Since Conjecture 1.7 is true, then $\mu_{i}$’s are indeed independent of $g(t)$ and thus well-defined for manifolds with unbounded curvature. Therefore we can remove the boundedness condition in Conjecture 1.7.
Corollary 1.10.
Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative curvature and Euclidean volume growth. Let $g(t)$ $(t\in[0,+\infty)$ be as complete solution Kähler-Ricci flow with initial metric $g$ as in Theorem 1.9, then
(1) $\mu_{i}(p)$ is independent of the choice of $p$ for any $1\leq i\leq n$.
(2) $\mu_{i}=d_{\operatorname{min}}^{(i)}-1$ for any $1\leq i\leq n$.
(3) $\operatorname{ACSD}(M,g(t))$ is invariant along $g(t)$.
Note that by Claim 4.2 in [16], the degrees of functions in $\mathcal{O}_{P}(M)$ also can be “simultaneously diagonalized” along the blowdown sequence, for details see Theorem 2.4 (5) in Section 2. As a by-product of Corollary 1.10, we can describe the Lyapunov regularity in (3) of Theorem 1.6 using the polynomial ring $\mathcal{O}_{P}(M)$. In some sense, it bridges the two distinct proofs of Yau’s uniformization conjecture in the case of maximal volume growth.
Corollary 1.11.
In (3) of Theorem 1.6, the basis of $V_{i}$ can be taken as $\left\{\frac{\partial}{\partial f_{i1}}(x),\cdots,\frac{\partial}{\partial f_{im_{i}}}(x)\right\}$. Here $\operatorname{dim}V_{i}=m_{i}$ is the multiplicity of $\mu_{i}$, $\sum\limits_{i=1}^{l}m_{i}=n$, and
$f_{is}\in\mathcal{O}_{P}(M)\text{ with }\operatorname{deg}(f_{is})=d_{\operatorname{min}}^{(i)}$
for any $1\leq s\leq m_{i}$.
Moreover, $\{f_{11},\cdots,f_{1m_{1}},\cdots,f_{l1},\cdots,f_{lm_{l}}\}$ can serve as a biholomorphism from $M$ onto $\mathbb{C}^{n}$.
This paper is organized as follows. Section 2 is some basic preliminary results and some simple conclusions that will be used later. In section 3, we will prove Theorem B by noting that any tangent cone of $M$ also arises as a tangent cone of corresponding expanding Kähler-Ricci solitons in Theorem 1.6. Also we prove Corollary 1.11. Section 4 contains the proof of Theorem A.
Acknowledgements
I would like to express my sincere gratitude to my advisor, Professor Yihu Yang, for providing constant support. I would like to extend my thanks to Professor Gang Liu for many valuable discussions.
2. Preliminary results
2.1. Structure of Kähler manifolds with maximal volume growth and the tangent cones
We first recall the main result in [16], and prove some relevant facts that will be used later.
Theorem 2.1.
Let $(M^{n},g,p)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth. Let $\operatorname{AVR}(M,g)=\nu>0$. Then
(1) $M$ is biholomorphic to $\mathbb{C}^{n}$.
(2) There exists an n-tuple polynomial growth holomorphic functions $(f_{1},\cdots,f_{n})$ serving as a proper biholomorphism onto $\mathbb{C}^{n}$. These functions satisfy
$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p,1)}f_{i}\overline{f_{j}}=\delta_{ij},\quad f_{i}(p)=0\quad\text{ for any }i,j\in\{1,\cdots,n\}.$ (2.1)
Moreover, any element in $\mathcal{O}_{P}(M)$ is a polynomial of these functions, meaning that $\mathcal{O}_{P}(M)\cong\mathbb{C}[f_{1},\cdots,f_{n}]$.
(3) Any two such global coordinates (consisting of n polynomial growth holomorphic functions satisfying (2.1), serving as a proper biholomorphism onto $\mathbb{C}^{n}$, and generating $\mathcal{O}_{P}(M)$) differ by a constant orthogonal transformation.
Proof.
The proof of these conclusions is implicitly contained in [11, 16]. For reader’s convenience, we give a thorough explanation. (1) and (2) are contained in Section 4 in [16].
The properness is also contained in [11]. It was shown that there exists a constant $D=D(n,\nu)$ such that let $(g_{1},\cdots,g_{k})$ be a linearly independent basis of $\mathcal{O}_{P}(M)$, where $k=\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{P}(M)-1$, such that
$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p,1)}g_{i}\overline{g_{j}}=\delta_{ij},\quad g_{i}(p)=0\quad\text{ for any }i,j\in\{1,\cdots,k\}.$
and
$\min\limits_{\partial B(p,r)}\sum\limits_{i=1}^{k}|g_{i}|^{2}\geq c(n,\nu)r^{2}$
holds for any $r>0$ and some constant $c=c(n,\nu)$.
Moreover, $(g_{1},\cdots,g_{k})$ serves as an embedding of $M$ onto an affine variety in $\mathbb{C}^{k}$.
Take an $n$-tuple algebraically independent holomorphic functions with minimal degrees in $(g_{1},\cdots,g_{k})$, denoted by $(f_{1},\cdots,f_{n})$. As in section 4 in [16], $(f_{1},\cdots,f_{n})$ serves as a biholomorphism of $M$ onto $\mathbb{C}^{n}$, and any polynomial growth function is a polynomial of them.
Since any $g_{i}$ is a polynomial of these $(f_{1},\cdots,f_{n})$ with degree no larger than $D$ by the three-circle theorem, by a contradiction argument we get
$\min\limits_{\partial B(p,r)}\sum\limits_{i=1}^{n}|f_{i}|^{2}\geq c(n,\nu)r^{\frac{2}{D}}$ (2.2)
holds for any $r>0$ and some constant $c=c(n,\nu)$. In other words, $(f_{1},\cdots,f_{n})$ is proper.
For (3), given any two such global coordinates $(f_{1},\cdots,f_{n})$ and $(h_{1},\cdots,h_{n})$. Assume their degrees are arranged in the nondecreasing order, then the matrix $\left(\frac{\partial h_{i}}{\partial f_{j}}\right)$ and $\left(\frac{\partial f_{i}}{\partial h_{j}}\right)$ are both polynomial matrices of $(f_{1},\cdots,f_{n})$ and are inverse of each other. Therefore they must both be constant matrices. Then due to the normalization condition (2.1), these two matrices are both orthogonal. ∎
In fact, the minimal degrees of all coordinates in Theorem 2.1 (3) described above are just the refined minimal order $\overrightarrow{d_{\operatorname{min}}}$ defined in the introduction, and it has been shown in [16].
Corollary 2.2.
Conjecture 1.5 (1) holds.
Proof.
If $M$ is of nonmaximal volume growth. By Theorem 2 in [12], $\mathcal{O}_{P}(M)=\mathbb{C}$, i.e. $\overrightarrow{d_{\operatorname{min}}}=\overrightarrow{+\infty}$, which is independent of the choice of $p$.
If $M$ is of maximal volume growth, due to [16], there exists a strictly increasing sequence $1\leq d_{1}<d_{2}<\cdots$, so that $\mathcal{O}_{P}(M)=\mathbb{C}\oplus\left(\mathcal{O}_{d_{1}}(M)/\mathbb{C}\right)\oplus\left(\mathcal{O}_{d_{2}}(M)/\mathcal{O}_{d_{1}}(M)\right)\oplus\cdots$ as a complex vector space. For any $k\in\mathbb{N}$, pick a maximal linearly independent vectors $f_{k1},\cdots,f_{km_{k}}$ of $\mathcal{O}_{d_{k}}(M)$ so that they form a basis of $\mathcal{O}_{d_{k}}(M)/\mathcal{O}_{d_{k-1}}(M)$ as quotient of vector spaces and no element in the span is given by polynomials of $\mathcal{O}_{d_{k-1}}(M)$. Then the first $n$ functions $\{f_{11},\cdots,f_{1m_{1}},f_{21},\cdots\}$ form a global coordinate in Theorem 2.1 after normalization as (2.1).
[…]
Remark 2.3.
(1) As can be seen in the proof, if $M$ is of maximal volume growth, Corollary 2.2 still holds without the nonsplitting condition on the universal cover.
(2) For simplicity, if $M$ is of maximal volume growth we just call the coordinate as in (2) in Theorem 2.1 a canonical coordinate on $M$, with degrees $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$. And we arrange them in non-decreasing order of their degrees.
Let $M_{\infty}$ be a tangent cone of infinity of $M$, whose structure has been studied in [14, 15, 9, 19, 18]. Let’s summarize these results in the following:
Theorem 2.4.
Let $(M,g)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth. Let $(M_{\infty},p_{\infty},d_{\infty})$ be a tangent cone at infinity of $M$, then:
(1) $(M_{\infty},p_{\infty},d_{\infty})$ is a complex manifold which is biholomorphic to $\mathbb{C}^{n}$. Moreover, it’s a metric Kähler cone with $BK\geq 0$” (see Section 5 in [19] for relevant definitions).
(2) There exists a complete Kähler-Ricci flow solution $h(t)$ on $M_{\infty}$ for $t\in[0,+\infty)$ with nonnegative bisectional curvature such that $(M_{\infty},p_{\infty},d_{\infty})$ is also the Gromov-Hausdorff limit of $h(t)$ as $t\rightarrow 0$. Moreover, $(M_{\infty},h(t))$ is a gradient Kähler-Ricci soliton with nonnegative bisectional curvature.
(3) $\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M_{\infty})=\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)$ for any $d\geq 0$. And there exists a n-tuple of polynomial growth functions $(z_{1\infty},\cdots,z_{n\infty})$ serving as a proper biholomorphism onto $\mathbb{C}^{n}$. Each $z_{i\infty}$ is of homogeneous degree $d_{\operatorname{min}}^{(i)}$, i.e. there exists a eigenfunction $\phi_{i}$ of the Laplacian operator on the link such that $z_{i\infty}=r^{d_{\operatorname{min}}^{(i)}}\phi_{i}$ on $M_{\infty}$. They satisfy
$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p_{\infty},1)}z_{i\infty}\overline{z_{j\infty}}=\delta_{ij},\quad z_{i\infty}(p_{\infty})=0\quad\text{ for any }i,j\in\{1,\cdots,n\}.$ (2.3)
Moreover, any element in $\mathcal{O}_{P}(M_{\infty})$ is a polynomial of these functions. In particular, $\mathcal{O}_{P}(M_{\infty})\cong\mathbb{C}[z_{1\infty},\cdots,z_{n\infty}]$.
(4) Any two such global coordinates (consisting of n polynomial growth homothetically homogeneous, holomorphic functions satisfying (2.3), serving as a proper biholomorphism onto $\mathbb{C}^{n}$, and generating $\mathcal{O}_{P}(M_{\infty})$) differ by a constant orthogonal transformation.
(5) Suppose there is a blow-down sequence $(M_{i},g_{i},p_{i})=(M,r_{i}^{-2}g,p)$ converges to $(M_{\infty},d_{\infty},p_{\infty})$ in the Gromov-Hausdorff sense. Let $(z_{1},\cdots,z_{n})$ be a canonical coordinate on $M$. Normalize them to $(z_{1k},\cdots,z_{nk})$ on $M_{k}$ such that
$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p_{k},1)}z_{ik}\overline{z_{jk}}=\delta_{ij},\quad z_{ik}(p_{k})=0\quad\text{ for any }i,j\in\{1,\cdots,n\},k\in\mathbb{N}.$ (2.4)
Then $(z_{1k},\cdots,z_{1n})$ sub-sequentially converges to a global coordinate $(z_{1\infty},\cdots,z_{n\infty})$ as in (3) on $M_{\infty}$ uniformly on compact subsets as $k\rightarrow+\infty$.
(6) The volume of the unit ball in $M_{\infty}$ is $\frac{\omega_{2n}}{d_{\operatorname{min}}^{(1)}\cdots d_{\operatorname{min}}^{(n)}}$. In particular, Conjecture 1.5 (2) holds.
Proof.
(1) and (2) were proved in Proposition 6.1 in [19].
For (3), for any $d\geq 0$, $\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M_{\infty})=\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)$ by Proposition 4.1 in [16]. Now we show the existence of such coordinate.
Firstly we claim that the homothetic vector field $r\frac{\partial}{\partial r}$ is holomorphic on $M_{\infty}$. By Theorem 1.9 there exists a long-time Kähler-Ricci flow solution $g(t)$ on $M$ such that $(M_{\infty},J_{\infty},p_{\infty},h(t))$ is the pointed Cheeger-Hamilton limit of a blow-down sequence of $g(t)$, i.e. $(M,J,p,g_{i}(t)=\frac{1}{t_{i}}g(t_{i}t)$ for some sequence $t_{i}\rightarrow+\infty$. In other words, $J$ converges to $J_{\infty}$ through a sequence of diffeomorphisms exhausted on $M_{\infty}$. By Claim 4.3 in [16], the complexification of $r\frac{\partial}{\partial r}$ is in the span of several holomorphic vector fields on $M_{\infty}$. Moreover, $(r\frac{\partial}{\partial r})\mathcal{O}_{d}(M_{\infty})\subset\mathcal{O}_{d}(M_{\infty})$ holds for any $d\geq 0$. The claim is confirmed.
Therefore, $r\frac{\partial}{\partial r}$ is a contracted holomorphic field on $M_{\infty}$, from a result in [25], $M_{\infty}$ is biholomorphic to $\mathbb{C}^{n}$. Now follow Section 4 in [16], replace the holomorphic vector field $X$ in the proof by $r\frac{\partial}{\partial r}$, we conclude the proof of (3) except the homogenity of these coordinates. This is also obvious since we can just choose the highest order terms of each coordinate functions and they certainly form a new biholomorphism onto $\mathbb{C}^{n}$.
The proof of (4) is the same as the proof of (3) in Theorem 2.1.
For (5), the mean value inequality [10] implies that,
$M_{z_{ik}}(\frac{1}{2})=\sup\limits_{B(p_{k},\frac{1}{2})}|z_{ik}|\leq C(n),$
holds for any $1\leq i\leq n$ and $k\in\mathbb{N}$. By the three-circle theorem in [13] and Cheng-Yau gradient estimate [5],
$M_{z_{ik}}(r)\leq C(n)r^{d_{\operatorname{min}}^{(i)}},\quad|dz_{ik}|\leq C(n)r^{d_{\operatorname{min}}^{(i)}-1}$
on $B_{p_{i}}(r)$ for any $1\leq i\leq n$, $k\in\mathbb{N}$ and $r\geq\frac{1}{2}$. Arzela-Ascoli theorem implies that $z_{ik}$ converges to a holomorphic function $z_{i\infty}$ of homogeneous degree no more than $d_{\operatorname{min}}^{(i)}$ uniformly on compact subsets. By (2.2) and the Gromov-Hausdorff convergence,
$\min\limits_{\partial B(p_{\infty},r)}\sum\limits_{i=1}^{n}|z_{i\infty}|^{2}\geq c(n,\nu)r^{\frac{2}{D(n,\nu)}}>0.$ (2.5)
Firstly, we show these functions are algebraically independent. Suppose not, then there exists a nontrivial polynomial $P\in\mathbb{C}[x_{1},\cdots,x_{n}]$ such that
$P(z_{1\infty},\cdots,z_{n\infty})\equiv 0$
on $M_{\infty}$. Lift to $M_{k}$ for large $k$ and use the properness (2.4) of $(z_{1k},\cdots,z_{nk})$, it implies $|P(z_{1k},\cdots,z_{nk})|$ can be arbitrarily small on arbitrarily large ball in $\mathbb{C}^{n}$. Thus $P\equiv 0$. And $z_{i\infty}$ must of homogeneous degree $d_{\text{min}}^{(i)}$ by the dimension estimate in (3), because on $M_{\infty}$ the minimal degree of a n-tuple algebraically independent holomorphic homogeneous functions has to be $\overrightarrow{d_{\operatorname{min}}}$. Therefore, (4) implies that $(z_{1\infty},\cdots,z_{n\infty})$ must form a global coordinate as in (3).
For (6), the first statement was proved in the proof of Theorem 3 in page 13 in [18]. This concludes the proof of Conjecture 1.5 (2) in the maximal volume growth case. The case of nonmaximal volume growth is trivial. ∎
Remark 2.5.
(1) As can be seen in the proof, if $M$ is of maximal volume growth, Conjecture 1.5 (2) still holds without the nonsplitting condition on the universal cover.
(2) For simplicity, we just call the coordinate as in (2) in Theorem 2.4 a canonical coordinate on $M_{\infty}$, with degrees $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$. And we arrange them in non-decreasing order of their degrees.
A natural corollary is the following homogeneity property of $\mathcal{O}_{P}(M)$:
Corollary 2.6.
Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth. Then for any $f,g\in\mathcal{O}_{P}(M)$,
$\operatorname{deg}(fg)=\operatorname{deg}(f)+\operatorname{deg}(g).$
Proof.
Take a canonical coordinate $(f_{1},\cdots,f_{n})$ on $M$ with
$f_{i}(p)=0\quad\text{ for any }i\in\{1,\cdots,n\}.$
For any sequence $r_{k}\rightarrow+\infty$, on $(M_{k},p_{k})=(r^{-2}_{k}M,p)$, we can normalize them such that
$f_{ik}=\frac{f_{i}}{M_{ik}},\quad\text{ where }M_{ik}=\sup\limits_{B_{g}(p,r_{k})}|f_{i}|$
for any $i\in\{1,\cdots,n\}$ and $k\in\mathbb{N}$. Therefore on each $M_{k}$,
$\sup\limits_{B(p_{k},1)}|f_{ik}|=1.$
Suppose $(M_{i},p_{i})\rightarrow(M_{\infty},p_{\infty})$ in the sense of Gromov-Hausdorff convergence, as in Proposition 5 in [15], each $f_{ik}\rightarrow f_{i\infty}$ with $f_{i\infty}$ is of homoegeneous degree $d_{\operatorname{min}}^{(i)}$.
As these functions generate $\mathcal{O}_{P}(M)$, we just need to prove: for any $f=f_{1}^{k_{1}}\cdots f_{n}^{k_{n}}$, $\operatorname{deg}(f)=\sum\limits_{i=1}^{n}k_{i}d_{\operatorname{min}}^{(i)}$.
Now let $f=f_{1}^{k_{1}}\cdots f_{n}^{k_{n}}$, by the three-circle theorem, for any $a>1$
$\lim_{r\to\infty}\frac{M_{f}(ar)}{M_{f}(r)}=a^{\operatorname{deg}(f)}.$
On the other hand,
$\lim_{r\to\infty}\frac{M_{f}(ar)}{M_{f}(r)}$ $=\lim\limits_{r\to\infty}\frac{M_{f_{1}^{k_{1}}\cdots f_{n}^{k_{n}}}(ar)}{M_{f_{1}^{k_{1}}\cdots f_{n}^{k_{n}}}(r)}$
$=\lim\limits_{k\to\infty}\frac{M_{f_{1}^{k_{1}}\cdots f_{n}^{k_{n}}}(ar_{k})}{M_{f_{1}^{k_{1}}\cdots f_{n}^{k_{n}}}(r_{k})}$
$=\lim\limits_{k\to\infty}\frac{M_{f_{1k}^{k_{1}}\cdots f_{nk}^{k_{n}}}(a)}{M_{f_{1k}^{k_{1}}\cdots f_{nk}^{k_{n}}}(1)}$
$=\frac{M_{f_{1\infty}^{k_{1}}\cdots f_{n\infty}^{k_{n}}}(a)}{M_{f_{1\infty}^{k_{1}}\cdots f_{n\infty}^{k_{n}}}(1)}$
$=a^{\sum\limits_{i=1}^{n}k_{i}d_{\operatorname{min}}^{(i)}}$
Compare these we complete the proof. ∎
Recently, Chu-Hao [7] obtain an optimal rigidity of the dimension estimate for polynomial growth holomorphic functions. Now we give an alternative proof, which seems much simpler.
Theorem 2.7.
Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature. Suppose there exists $d\geq 1$ such that
$\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)>\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(\mathbb{C}^{n})-\binom{n+d-2}{d-1}$,
then $M$ is biholomorphically isometric to $\mathbb{C}^{n}$ with Euclidean metric.
Proof.
Suppose $M$ is of maximal volume growth, then take $(f_{1},\cdots,f_{n})$ be a canonical coordinate with minimal degree $(d^{(1)}_{\operatorname{min}},\cdots,d^{(n)}_{\operatorname{min}})$ with $d^{(i)}_{\operatorname{min}}\geq 1$ in non-decreasing order.
From Corollary 2.6, $\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)=\#\{(m_{1},\cdots,m_{n})\in\mathbb{Z}^{n}_{\geq 0}\mid\sum\limits_{i=1}^{n}m_{i}d^{(i)}_{\min}\leq d\}$. We know that $M$ is biholomorphically isometric to $\mathbb{C}^{n}$ if and only if $d^{(i)}_{\min}=1$ for all $i$. Suppose $M$ is not isometric to $\mathbb{C}^{n}$. Then $d^{(n)}_{\min}>1$, then
$\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)$ $=\#\{(m_{1},\cdots,m_{n})\in\mathbb{Z}^{n}_{\geq 0}\mid\sum\limits_{i=1}^{n}m_{i}d^{(i)}_{\min}\leq d\}$
$\leq\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(\mathbb{C}^{n})-\#\{(m_{1},\cdots,m_{n-1})\in\mathbb{Z}^{n-1}_{\geq 0}\mid\sum\limits_{i=1}^{n}m_{i}<d\}$
$=\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(\mathbb{C}^{n})-\binom{n+d-2}{d-1}$.
with equality holds if and only $d^{(i)}_{\operatorname{min}}=1$ for $i\leq n-1$ and $1<d^{(n)}_{\operatorname{min}}<\frac{d}{d-1}$. This contradicts with our assumption.
Now suppose the universal cover $\widetilde{M}$ splits. By the proof of Corollary 7.1 in [11], $\widetilde{M}=M_{1}^{m_{1}}\times M_{2}^{n-m_{1}}$, where $M_{1},M_{2}$ are both simply connected and has nonnegative bisectional curvature. Also $M_{1}$ is of maximal volume growth. $\mathcal{O}_{d}(M)$ can be seen as the $G$-invariant subring of $\mathcal{O}_{d}(M_{1})$, $G$ is a compact Lie group acting on $M_{1}$.
Therefore
$\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(\mathbb{C}^{n})-\binom{n+d-2}{d-1}<\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)\leq\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M_{1}),$
From the paragraph above, $M_{1}$ must be biholomorphically isometric to $\mathbb{C}^{n}$, i.e. $\widetilde{M}=\mathbb{C}^{n}$. And the $G$-action on $\mathbb{C}^{n}$ must be the deck transformation of $\pi_{1}(M)$.
Claim: Let $\pi:\widetilde{M}\rightarrow M$ be the covering map, $f\in\mathcal{O}_{P}(M)$, then $\operatorname{deg}(f)=\operatorname{deg}(f\circ\pi)$.
It’s obvious since $\pi(B(\tilde{p},r))=B(p,r)$ for any $\pi(\tilde{p})=p$ and any $r>0$.
Since $\mathcal{O}_{P}(M)$ is finitely generated, take a family of minimal generators of it, say $\{f_{1},\cdots,f_{N}\}\subset\mathcal{O}_{P}(\mathbb{C}^{n})$. By the claim we know that $\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)=\#\{\text{monomial of }f_{1},\cdots,f_{N}\text{ which is in }\mathcal{O}_{d}(\mathbb{C}^{n})\}$.
Similar argument as above shows that
$\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)\leq\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(\mathbb{C}^{n})-\binom{n+d-2}{d-1}$
unless $\{f_{1},\cdots,f_{N}\}=A\{z_{1},\cdots,z_{n}\}+b$ for some constant $A,b$, in which case $\pi_{1}(M)$ is trivial. Therefore we complete the proof. ∎
Remark 2.8.
Similar argument can also prove the splitting theorem in Section 3 in [7], by using Theorem 4.1 in [23].
2.2. Gradient expanding Kähler-Ricci solitons
Definition 2.9.
A Kähler-Ricci soliton consists of a triple $(M,g,X)$, where $M$ is a Kähler manifold, $X$ is a holomorphic vector field on $M$, and $g$ is a complete Kähler metric on $M$ whose Kähler form $\omega$ satisfies
$\operatorname{Ric}-\frac{1}{2}\mathcal{L}_{X}\omega+\lambda\omega=0$
for some $\lambda\in\{-1,0,1\}$. A Kähler-Ricci soliton is said to be expanding if $\lambda=1$. Here $X$ is called the soliton vector field. If in addition, $X=\nabla^{g}f$ for some real-valued smooth function $f$ on $M$, then we say $(M,g,X)$ is a gradient Kähler-Ricci soliton. In this case, we call $f$ the potential function of the soliton. It’s equivalent to say
$\operatorname{Ric}_{i\bar{j}}-f_{i\bar{j}}+\lambda g_{i\bar{j}}=0,\quad f_{ij}=f_{\bar{i}\bar{j}}=0.$
They we list some important properties of gradient expanding Kähler-Ricci solitons with nonnegative Ricci curvature which can be found in [2, 6, 30, 4, 28].
Proposition 2.10.
Let $(N^{n},J,g,f)$ be a complete noncompact gradient expanding Kähler-Ricci solitons with nonnegative Ricci curvature. Then
(1) $R+|\nabla^{g}f|^{2}-f$ is constant on $N$, therefore we always normalize $f$ so that $f=R+|\nabla^{g}f|^{2}$.
(2) The potential $f$ is a strictly convex exhaustion function with $\nabla^{g}f$ has the unique zero at $O\in N$, and $R$ attains its maximum at $O$. In particular, if $N$ has nonnegative bisectional curvature, then $N$ has bounded curvature.
(3) $(N,g)$ is biholomorphic to $\mathbb{C}^{n}$ and has maximal volume growth.
(4) $\nabla^{g}f$ is a complete vector field. Let $\varphi(t)$ be a family of biholomorphism of $(N,J)$ with
$\frac{\partial}{\partial t}\varphi(t-1)(x)=-\frac{1}{t}\nabla^{g}f(\varphi(t-1)(x)),\varphi(0)=\text{id}.$ (2.6)
It follows that
$g(t)=t\varphi(t-1)^{*}g$ (2.7)
solves the Kähler-Ricci flow equation on $[0,+\infty)$ with $g(1)=g$.
(5) ((13) in [28]) The potential function $f$ satisfies the estimate
$f(\varphi(O,t))\leq f(O)e^{2t}.$ (2.8)
3. Proof of Theorem B
The key observation is the following which may be well-known for experts in the fields. However, for reader’s convenience, we give a proof here.
Lemma 3.1.
Let $(N,J,h,O,f)$ be a complete noncompact expanding Kähler-Ricci soliton with nonnegative bisectional curvature and we normalize it by assuming $f=R+|\nabla^{h(1)}f|^{2}$. Let
$h(t)=t\varphi(t-1)^{*}h$
be the unique Kähler-Ricci flow solution on $[0,+\infty)$ such that $h(1)=h$. Where
$\frac{\partial}{\partial t}\varphi(t-1)(x)=-\frac{1}{t}\nabla^{g}f(\varphi(t-1)(x)),\varphi(0)=\text{id}.$
Then $(N,h(1))$ has a unique tangent cone at infinity $(N_{\infty},h_{\infty},O_{\infty})$. Moreover, $(N,h(t),O)\rightarrow(N_{\infty},h_{\infty},O_{\infty})$ as $t\rightarrow 0$ in the Gromov-Hausdorff sense.
Proof.
Replace $(M,g)$ in Theorem 1.6 by $(N,h(1))$, we conclude that there exists a metric Kähler cone $N_{\infty}$ such that $(N,h(t),O)\rightarrow(N_{\infty},h_{\infty},O_{\infty})$ as $t\rightarrow 0$ in the Gromov-Hausdorff sense. So it remains to prove this cone $N_{\infty}$ is the unique tangent cone of $(N,h(1))$. In fact we have
$(N_{\infty},h_{\infty},O_{\infty})$ $=\lim\limits_{t\to 0}(N,h(t),O)$
$=\lim\limits_{t\to 0}(N,t\varphi(t-1)^{\star}h(1),O)$
$=\lim\limits_{t\to 0}(N,th(1),\varphi(t-1)(O))$
$=\lim\limits_{t\to 0}(N,th(1),(O)).$
All the limits above are in the sense of Gromov-Hausdorff convergence. The second-to-last equality holds by the following claim:
Claim 1.
$\lim\limits_{t\to 0}d_{th(1)}(\varphi(t-1)(O),O)=0$.
Proof.
For any $s\leq 1$, by (2) and (2.8) in Proposition 2.10,
$|\nabla^{h(1)}f|^{2}\left(\varphi(s)(O)\right)\leq R+|f(\varphi(s)(O))|\leq C(1+e^{2})\leq C.$
Integrating along $\varphi(s)(O)$, we obtain
$\sqrt{t}d_{h(1)}(\varphi(t-1)(O),O)\leq\sqrt{t}\int_{t}^{1}\frac{C}{s}ds=-C\sqrt{t}\log(t)\rightarrow 0\text{ as }t\rightarrow 0.$
Thus, we now have a clear geometric framework characterizing the long-time asymptotics of the Kähler-Ricci flow on a complete noncompact Kähler manifold with bounded nonnegative bisectional curvature and maximal volume growth, along with its relationship to its tangent cones.
Now we prove Theorem B.
Proof of Theorem B:
Following the notation of Theorem 1.6, take the Poincaré coordinate $(w_{1},\cdots,w_{n})$ as in page 2622 in [28] on $(N,h(1),O)$ and normalize them such that
$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B_{h(1)}(O,1)}{w_{i}}\overline{w_{j}}=\delta_{ij},\quad w_{i}(O)=0\quad\text{ for any }i,j\in\{1,\cdots,n\}.$
According to the proof of Theorem 1.4 in [28], the aforementioned Poincaré coordinate $(w_{1},\cdots,w_{n})$ is a canonical coordinate on $(N,h(1),O)$. Following the procedure of (5) in Theorem 2.4 on $(N,h(1),O)$ and $(M,g,p)$, we get two canonical coordinates $(z_{1\infty},\cdots,z_{n\infty})$ and $(w_{1\infty},\cdots,w_{n\infty})$ on $N_{\infty}=M_{\infty}$, with degree $\overrightarrow{d_{\operatorname{min}}}(M)$ and $\overrightarrow{d_{\operatorname{min}}}(N)$. Thus from (4) in Theorem 2.4,
$\overrightarrow{d_{\operatorname{min}}}(M)=\overrightarrow{d_{\operatorname{min}}}(N)=\{\mu_{1}+1,\cdots,\mu_{n}+1\}.$
Therefore, (1),(2) has been proved. For (3), note that in [18], Liu has proved that $\operatorname{ASCD}(M,g(t))=4n(\sum\limits_{i=1}^{n}d^{(i)}_{\operatorname{min}}-n)=4n(\sum\limits_{i=1}^{n}\mu_{i})$ is invariant along $t$. The proof is complete.
As an application, we can use it to prove Corollary 1.11:
Proof of Corollary 1.11:
Firstly, we assume $(M,g)$ has bounded curvature. Take a canonical coordinate $(f_{1},\cdots,f_{n})$ on $(M,g)$, with degree $\overrightarrow{d_{\operatorname{min}}}$. By comparing the distance function $d_{g(t)}(x,p)$ along the flow using the shrinking balls lemma by Simon and Topping (Corollary 3.3 in [26]), we may conclude that $(f_{1},\cdots,f_{n})$ is still a global coordinate on $(M,g(t))$, with the same degree $\overrightarrow{d_{\operatorname{min}}}$. Let $g_{k}(t)=\frac{1}{t_{i}}g(t_{i}t)$, we normalized them such that
$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B_{g_{k}(1)}(O,1)}{f_{ik}}\overline{f_{jk}}=\delta_{ij},\quad f_{ik}(p)=0\quad\text{ for any }i,j\in\{1,\cdots,n\},k\in\mathbb{N}.$
which makes $(f_{1k},\cdots,f_{nk})$ serving as a canonical coordinate on $(M,g_{i}(1))$.
Following the same proof as in (5) in Theorem 2.4, these $(f_{1k},\cdots,f_{nk})$ converges smoothly to $(f_{1\infty},\cdots,f_{n\infty})$ on $(N,h(1))$ and serving as a canonical coordinate on $N$. Due to (3) in Theorem 2.1, WLOG we can assume $(f_{1\infty},\cdots,f_{n\infty})$ is just a Poincaré coordinate on $(N,h(1))$. Therefore
$\operatorname{Ric}_{h(1)}\left(\frac{\frac{\partial}{\partial f_{i\infty}}}{|\frac{\partial}{\partial f_{i\infty}}|},\overline{\frac{\frac{\partial}{\partial f_{i\infty}}}{|\frac{\partial}{\partial f_{i\infty}}|}}\right)=\mu_{i}.$
Finally, by the smooth convergence of $g_{i}(1)$ and the proof of Theorem 1.2 in [3], the Corollary is proved. In fact, for any $i$ and any nonzero $v\in V_{1}\oplus\cdots\oplus V_{i}$ but $v\notin V_{1}\oplus\cdots\oplus V_{i-1}$,
$\lim\limits_{t\to\infty}t\operatorname{Ric}(v(t),\overline{v(t)})=\mu_{i}$
always holds.
If $(M,g)$ has unbounded curvature, thanks to Theorem 1.9, above argument can still be applied.
4. Proof of Theorem A
Proof of Theorem A:
Firstly we assume $M$ is of nonmaximal volume growth, assume there exists a nonzero polynomial holomorphic $n$-form $s\in H^{0}_{P}(M,\mathcal{K}_{M})$, the Poincaré-Lelong equation states that
$\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\log\|s\|^{2}=(s)+\operatorname{Ric}\geq 0,$
where $(s)$ is the zero divisor of $s$. In other words, $h=\log(\|s\|^{2}+1)$ is a smooth plurisubharmonic function of logarithmic growth. Since the universal cover of $M$ does not split, the heat-flow method of Ni-Tam [23] can be applied to perturb $h$ to $h^{\prime}$ which is a smooth plurisubharmonic function of still logarithmic growth (for details, see Theorem 0.1 (iii) in [23]). Then we solve the $\bar{\partial}$-equation by applying the Hörmander $L^{2}$-estimate similarly as the proof of Lemma 3 in [12], there exists a nonconstant polynomial growth holomorphic function on $M$, contradicting with Theorem 2 in [12]. As a result, $D_{\operatorname{min}}=+\infty=\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n$.
Now we assume $M$ is of maximal volume growth. Let $(z_{1},\cdots,z_{n})$ be a canonical coordinate of $M$ with degrees $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$ as stated in Remark 2.3. By gradient estimate [5],
$\operatorname{deg}(dz_{1}\wedge\cdots\wedge dz_{n})\leq\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n,$
therefore $D_{\operatorname{min}}\leq\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n$.
The proof of the opposite direction is similar to the proof of Corollary 1.5 in [28], but a little more complicated. Let $s\in H^{0}_{D_{\operatorname{min}}}(M,\mathcal{K}_{M})$, then note that Corollary 3 in [13] also holds for holomorphic sections of holomorphic line bundle with nonpoaitive curvature, there exists a constant $C>0$ such that
$M_{s}(r,0)\leq Cr^{D_{\operatorname{min}}}\quad\text{ for any }r>0.$ (4.1)
Fix $t_{k}\rightarrow+\infty$, there exists a long-time Kähler-Ricci flow solution $g(t)$ which satisfies properties in Theorem 1.9. And set $g_{k}(t)=\frac{1}{t_{i}}g(t_{i}t)$, assume $t_{k}\geq 1$.
Claim 2.
[…]
Proof.
By the shrinking balls lemma by Simon and Topping (Corollary 3.3 in [26]), for any $t_{k}>0$, we have
$d_{g}\leq d_{g(t_{k})}+C^{\prime}_{k}\sqrt{t_{k}}$
on $M\times[0,+\infty)$.
Also, thank to Corollary 1.10, (1.2) also holds for $M$ with unbounded curvature, which means that
$\lim\limits_{t\to\infty}\log\frac{\det(g_{i\bar{j}}(x,t))}{\det(g_{i\bar{j}}(x,0))}=-C.$
for some nonnegative constant $C$ which relies only on $M$ and is independent of $x$. Also by the Ricci flow equation,
$\frac{d}{dt}\log\frac{\det(g_{i\bar{j}}(x,t))}{\det(g_{i\bar{j}}(x,0))}=-R(x,t)\leq 0.$
So
$\log\frac{\det(g_{i\bar{j}}(x,t))}{\det(g_{i\bar{j}}(x,0))}\geq-C$
holds uniformly on $M\times[0,+\infty)$.
Therefore, on $(M,g(t_{k}))$, we have
$M_{s}(r,t_{k})\leq M_{s}(r+C^{\prime}_{k}\sqrt{t_{k}},0)\frac{\det(g_{i\bar{j}}(x,0))}{\det(g_{i\bar{j}}(x,t))}\leq C_{k}r^{D_{\operatorname{min}}}.$
The claim is proved. ∎
[…]
Thanks to Theorem 1.9, the boundedness condition of curvature in Theorem 1.6 can be removed. Using the notations in Theorem 1.6, recall the sequence $(U_{k},F_{k}^{*}J.F_{k}^{*}(g_{k}(1)),O)$ converges smoothly to $(N,J_{\infty},h(1),O)$ uniformly on compact subsets. Define
$s_{k}=\frac{s\mid_{F_{k}(U_{k})}}{\sup\limits_{B_{g_{k}(1)}(p,1)}\|s\|},\quad\tilde{s}_{k}=F_{k}^{*}s_{k}.$ (4.2)
Since $\mathcal{K}_{M}(g_{k}(1))$ has nonpositive curvature, three-circle theorem in [13] can be applied to get for any fixed $r>0$, for large $k$,
$\sup\limits_{B_{g_{k}(1)}(p,r)}\|s_{k}\|\leq r^{D_{\operatorname{min}}}.$
After pulling back by $F_{k}$, we obtain
$\sup\limits_{B_{F_{k}^{*}g_{k}(1)}(O,r)}\|s_{k}\|\leq r^{D_{\operatorname{min}}}.$ (4.3)
holds for large $k$.
Fix a nonvanishing holomorphic $n$-form $\Omega$ on $N$, then write $s_{k}=f_{k}\Omega_{k}$ on $U_{k}$, where $f_{k}$ is a holomorphic function with respect to $F_{k}^{*}J$ on $U_{k}$, $\Omega_{k}$ is a nonvanishing holomorphic $n$-form with respect to $F_{k}^{*}J$ and converges to $\Omega$ uniformly on each fixed compact subset in $N$ (For example, take $U_{k}$ be a sequence of exhausted Euclidean balls in $N\cong\mathbb{C}^{n}$).
Therefore there exists a constant $C=C(K)$ such that $\|\Omega_{k}\|_{F_{k}^{*}(g_{k}(1))}\geq C(K)$ uniformly on $K\subset N$ compact. Then we can deduces
$\sup\limits_{K}|f_{k}|\leq\frac{1}{C(K)}r^{D_{\operatorname{min}}}$
holds for all large $k$. Gradient estimate and Arzela-Ascoli theorem imply that $f_{k}\rightarrow f$ uniformly on compact subsets in $N$, and $f$ is holomorphic with respect to $J_{\infty}$. In other words, sub-sequentially $s_{k}$ converges to a holomorphic $n$-form $s^{\prime}=f\Omega$ on $N$.
From (4.2) and (4.3), $s^{\prime}$ is nonzero ($\sup\limits_{B_{h(1)}(O,1)}\|s^{\prime}\|=1$) and satisfies
$\sup\limits_{B_{h(1)}(O,r)}\|s_{k}\|\leq r^{D_{\operatorname{min}}}.$
Combining with (1.5) we obtain
$D_{\operatorname{min}}\geq D_{\operatorname{min}}(N)=\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}(N)-n=\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n.$
Therefore
$D_{\operatorname{min}}=\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n.$
A much simpler way is consider directly a tangent cone of $M$. Assume $(M,g_{i}=r_{i}^{-2}g)\rightarrow M_{\infty}$ in the Gromov-Hausdorff sense, where $r_{i}\rightarrow\infty$. Take a canonical coordinates $(z_{1},\cdots,z_{n})$ on $(M,g)$, we know that
$D_{\operatorname{min}}=\operatorname{deg}(dz_{1}\wedge\cdots\wedge dz_{n}).$
And from (4) we have known that
$D_{\operatorname{min}}\leq\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n.$
By the three-circle theorem, assume
$D_{\operatorname{min}}\leq\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n-\varepsilon$
for some $\varepsilon>0$.
Following the notations in Theorem 2.3 (5), and define
$\varphi_{i}=\log\|dz_{1i}\wedge\cdots\wedge dz_{ni}\|^{2}_{g_{i}}$
As in [18], since $\varphi_{i}$ are psh on $M_{i}$, by passing to a subsequence, we may assume $\varphi_{i}$ converges to a psh function on $M_{\infty}$ in $L_{\operatorname{loc}}^{1}$, denoted by $\varphi=\log|dz|^{2}$. Therefore $e^{\varphi_{i}}\rightarrow e^{\varphi}$ almost everywhere.
By the three circle theorem as in the proof of Theorem 2.3 (5), there exists a $C$ such that
$e^{\varphi_{i}}\leq Cr_{i}^{2D_{\operatorname{min}}}\leq Cr_{i}^{2(\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n-\varepsilon)}.$
uniformly. Thus $e^{\varphi}\leq Cr^{2(\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n-\varepsilon)}$ almost everywhere, which contradicts with Claim 1 in [18] if $\varepsilon>0$, since $e^{\varphi}$ is homogeneous of degree $2(\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n)$. This directly implies that $D_{\operatorname{min}}=\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n.$